OpenStudy (batman19991):

If a circle is enlarged to have a circumference that is four times the original, then the original radius was multiplied by four. True False

OpenStudy (anonymous):

\[2 \pi R=4*2 \pi r,R=4r\]

OpenStudy (batman19991):

what